# Novel View Synthesis for Cinematic Anatomy on Mobile and Immersive Displays

### v1

@@ 6 Conclusion
In this work we have demonstrated the use of differentiable 3D Gaussian splatting for novel view synthesis from path traced images of high resolution medical datasets. We have shown that the 3D Gaussian representation can be compressed — at hardly perceivable loss in image quality — to a size that enables download and storage on even mobile devices. Even though the Gaussian representation needs to be re-generated for every selected preset, even when many different presets are used the overall memory is still far below the memory required by the dataset. Computationally expensive path or ray tracing can be avoided at rendering time, enabling fast display on mid- and even low-end devices.
@@
We have also pointed at current limitations of 3DGS in the envisioned scenario. As the most crucial ones we see the current absence of support for clip planes and the quality degradation when semi-transparent volume rendering is performed. We have sketched future research directions to address these limitations, and we are confident that significant improvements in such scenarios can be achieved. Another challenging aspect that needs to be addressed in the future is the handling of time-varying dataset. We see more and more scanning technologies that can accurately measure blood flow and deforming tissue. Tailoring 3DGS for interactively monitoring and inspecting such dynamic processes is another important goal.
Acknowledgements.
HiP-CT data from the UCL led European Synchrotron Radiation Facility (ESRF) beamtime MD1290 (courtesy Lee, Walsh, Tafforeau et al.), Nature Methods, Walsh et al 2021[51].

### v2

@@ 6 Conclusion
We have demonstrated the use of differentiable 3DGS for novel view synthesis from path traced images of high resolution medical data sets. We have shown that the 3D Gaussian representation can be compressed — at hardly perceivable loss in image quality — to a size that enables download and storage on even mobile devices. The Gaussian representation needs to be re-generated for every selected preset, yet even for many presets the overall memory is still significantly below the memory required by the data set. Computationally expensive path tracing can be avoided at rendering time, enabling fast display on mid- and even low-end devices.
@@
We have also pointed at current limitations of 3DGS for CA. As the most important ones we see the current absence of support for clip planes and the quality degradation when a headlight is used. We have sketched future research directions to address these limitations, and we are confident that improvements can be achieved. There is also a pressing need to handle time-varying data sets, since we see more and more scanning technologies that can accurately measure blood flow and deforming tissue. Tailoring 3DGS for interactively visualizing such dynamic processes is another important goal.
Finally, we want to mention that besides CA we see in-situ visualization as another promising application of 3DGS. For data sets which are simulated on a supercomputer and are so large that they cannot be streamed out, images of the data set can be generated directly on the supercomputer and then streamed out to a system where novel view synthesis is performed. By using advanced implementations of 3DGS optimization, this might even become possible at rates enabling an explorative visual analysis.
Supplementary Material
A 3DGS Optimizations
A.1 Volume Guided Initialization
When using 3DGS, an initial set of 3D Gaussian kernels is first selected. These Gaussians are then removed, split or re-positioned, and the shape and appearance of the Gaussian kernels is optimized. Kerbl et al. [KKLD23] obtain the initial positions of the 3D Gaussians from the given images with structure from motion, or with random initialization where Gaussians are randomly positioned in the scene. For volume rendering, we randomly place Gaussians within the volume bounding box and set their initial color to grey. All other parameters are initialized as proposed by Kerbl et al. [KKLD23].
In Fig. 10 we exemplarily compare the effectiveness of the different initialization schemes based on optimization convergence for one of our test data sets. An initialization with the Gaussians’ positions and colors from a previous reconstruction is used as gold standard. As can be seen, while all initialization techniques reach the same level of fidelity, volume-guided initialization does so with less iteration steps. However, it is fair to say that in all of our experiments the performance improvements were overall not significant, so that we decided to use random initialization in all tests.
Figure 10: Test image reconstruction using different initialization schemes. Experiments were performed with Body.
Duration Size Points Train Images Test Images SSIM PSNR PSNR (Alpha)
Scene Preset Resolution
Brain 1 2k 108 Min 170 MB 5.1 M 87 12 0.64 20.73 32.76
2 2k 88 Min 156 MB 5.0 M 87 12 0.81 25.78 35.86
Kidney 1 2k 50 Min 73 MB 2.7 M 91 13 0.81 23.93 27.70
2 2k 49 Min 60 MB 2.2 M 87 12 0.87 27.89 32.85
Body 1 2k 58 Min 30 MB 1.0 M 87 12 0.89 28.93 31.54
2 2k 43 Min 28 MB 1.0 M 87 12 0.87 25.38 29.94
3 2k 43 Min 30 MB 1.0 M 87 12 0.88 27.65 30.40
4 1k 20 Min 55 MB 1.9 M 224 32 0.86 28.68 25.41
Table 4: High Quality reconstruction details for all scenes and presets.
Size Gaussians SSIM PSNR PSNR (Alpha)
Scene Preset Resolution
Brain 1 2k 69 4.8 M 0.63 20.67 32.48
2 2k 69 4.8 M 0.81 25.79 35.70
Kidney 1 2k 37 2.6 M 0.81 23.91 27.81
2 2k 30 2.1 M 0.87 27.70 32.58
Body 1 2k 7 1.0 M 0.88 28.51 30.60
2 2k 7 0.9 M 0.86 25.03 29.08
3 2k 7 1.0 M 0.87 27.17 29.03
4 1k 22 1.5 M 0.81 26.13 24.46
Table 5: High Range compression details for all scenes and presets
A.2 Mip Splatting
Scenes rendered with 3DGS can show severe artifacts when novel camera perspectives diverge from those the 3D Gaussian representation was optimized for. Yu et al. [YCH23] name the following two reasons for this behavior: Firstly, the 3D Gaussian representation exhibits frequencies that are too high to be faithfully reconstructed by the used sampling rate. Secondly, during splat-based rendering, a 2D dilation filter is applied that causes artefacts when zooming out and 2D splats become too small.
B Performance Statistics and Further Results
In Table 4 and Table 5, we provide detailed statistics using HR- and HQ-compression for all used presets. Additional qualitative results are shown in Fig. 13.
C View Selection
In Fig. 11, we compare the convergence rate of BOS-based view selection to the approach of Kopanas and Drettakis [KD23], using preset 4 of Body. The final energy term adapted from Kopanas and Drettakis [KD23] is plotted after selecting a total of 32 cameras using a variable number of tested candidate views. Fig. 12 shows the time it takes the view selection algorithms to select the 32 camera poses for these two strategies. BOS is able to find better maxima than pure random sampling of the candidate camera poses as described by Kopanas and Drettakis, albeit at a slightly higher computation time for the same number of generated candidate poses. The tests were run on a system with an AMD Ryzen 9 3900X 12-core (24-thread) CPU and an NVIDIA GeForce RTX 3090 GPU and averaged over 8 different random seeds. The BOS algorithm is run on the CPU, while the computation of the transmittance as described in the main manuscript is performed on the GPU using ray marching of a downscaled version of the volume.
Figure 11: Achieved energy (the higher the better) when 32 camera poses are selected to generate test images of Body. On the $x$-axis, the batch size of candidate poses in each iteration is shown. On the $y$-axis, the energy, as adapted from Kopanas and Drettakis [KD23], is shown.
Figure 12: Run time for selecting 32 camera poses for Body. On the $x$-axis, the batch size of candidate poses in each iteration is shown. On the $y$-axis, the run time is shown.